the pressure due to atmosphere is 1.013 into 10 to the power 5 pascals find the force exerted by the atmosphere on the top of a surface of a table 2 metres long and 1m wide

Grade:9

3 Answers

Arun
5 years ago

Well if you are talking about the earth's atmosphere then the standard value for earth's atmospheric pressure is 101.325 kPa or 1.01325 × 105 Pa

So 

F = P × Area of the table 

and Area of the table  = lb = 2 m2

Now by substituting the values we can get the desired answer

RAHUL ROHILLA
5 years ago
2.01X105you can use P=F/A formula for this
CHEETALI SHARMA
2 years ago
HERE THE ATMOSPHERIC PRESSURE IS GIVEN AS = 1 013×10^5 Pa
AREA OF THE TABLE IS GIVEN AS = 2m×1m = 2m ^2
NOW THE FORMULA FOR PRESSURE US FORCE ACTING PER UNIT AREA WHICH IS 
P=F/A
SO FORCE, F = P×A
WHICH IS,  F = 1.0132×10^5  ×  2 = 2.0264×10^5 Pa